OpenStudy (maddie82502):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (xmissalycatx):

First you will want to distribute the 5 to everything in the parentheses.. Your result would be (40 + 35) = 5x + 35 Next you would combine like terms on the left side which means you add 40 and 35 which is 75.. so 75 = 5x + 35 Next you would subtract 35 from both sides so that "5x" is by itself and you can solve for x.. so 75- 35= 40... therefore 40= 5x Now, you will divide both sides by 5, so that you get x. 40/5 is 8! x=8

OpenStudy (nethercreep333):

Let's solve your equation step-by-step. 5(8+7)=5x+(5)(7) Step 1: Simplify both sides of the equation. 75=5x+35 Step 2: Flip the equation. 5x+35=75 Step 3: Subtract 35 from both sides. 5x+35−35=75−35 5x=40 Step 4: Divide both sides by 5. 5x 5 = 40 5 x=8